OpenStudy (anonymous):

the lengths of the sides of a triangle are in the extended ratio 3:1:12. The perimeter is 4cm. What is the length of the longest side in centimeters? I have absolutely no idea, I need someone to show me how to do this.

OpenStudy (anonymous):

Is this it? 3/25 = x/400 ??

OpenStudy (anonymous):

So that would then be: 1200 = 25x 1200/25 = 25x/25 48 = X X = 48

OpenStudy (anonymous):

The length of the longest side = 48cm ??????